Using 10 ml of 12 M HCl stock solution, calculate the volume of water necessary to
create a 0.01 M solution of HCl. What is the final volume of the solution?
1
Expert's answer
2020-05-10T14:58:15-0400

For solutions (1 and 2) dilution the equation is:

C1*V1 = C2*V2, where C is concentration, V is volume

Final volume of the solution is:

V2 = C1*V1/C2 = 10 ml * 12 M / 0.01 M = 12000 ml = 12 L

To prepare is the volume of water needed is V(H2O) = 12000 ml (solution) - 10 ml (solution 12 M HCl) = 11990 ml

Answer: final solution volume is 12000 ml; volume of water needed to prepare it is 11990 ml